OpenStudy (anonymous):

Initial Value Problem: dA/ dt = 6A,,,,,,,,,,,,,, A(0) = 9, Solve for A(t) = ????

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

separate out the variables, keep terms with A on one side and t on other side of =

OpenStudy (anonymous):

I did that already, I got to the point where I have dA/dt = 6 dt after integrating you will get ln(a) = 6t. Am I right so far?

hartnn (hartnn):

add the constant of integration! ln A = 6t +C

hartnn (hartnn):

now you can use the fact that A(0) = 9 do you know what that means ?>

OpenStudy (anonymous):

I know but ln(0) is undefined,,,, this is where I got stuck.

hartnn (hartnn):

that means, its not clear to you :) A(0) = 9 means when \(\Large t=0, A=9\)

OpenStudy (anonymous):

Oh okay, so now we solve for C. Thanks a lot, I know that but had a long. Thanks again.

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

@hartnn One more question. After you solve for C, how do you seperate A as a function?

OpenStudy (anonymous):

ln(A) = 6t + ln(9)

hartnn (hartnn):

you want to isolate A ?

OpenStudy (anonymous):

I want A(t) = something.

hartnn (hartnn):

yep, that means you need to isolate A

hartnn (hartnn):

bring ln 9 on left side and use log property ln x - ln y = ln (x/y)

OpenStudy (anonymous):

ln(A/6) = 6t?

hartnn (hartnn):

ln(A/9) = 6t yes

hartnn (hartnn):

now again if you're good with logs, you shouldn't have any problems \(\Large \ln a=b \implies a=e^b\)

OpenStudy (anonymous):

so A/9 = e^(6t) ?

OpenStudy (anonymous):

@hartnn I got it thanks!!!!! I really appreciate your effort.

hartnn (hartnn):

welcome :)